OpenStudy (anonymous):

What is the cube root of 27x18? 3x6 3x9 9x6 9x9

OpenStudy (anonymous):

@Michele_Laino @Nnesha

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (michele_laino):

\[27*18=3^{3}*(2*3^{2})=3*2*3^{4}=6*(3^{2})^{2}=6*9^{2}\]

OpenStudy (anonymous):

@Michele_Laino Not an answer choice

OpenStudy (michele_laino):

ok! I write it better:

OpenStudy (michele_laino):

\[27*18=3^{3}*(2*3^{2})=3^{3}*18\]

OpenStudy (anonymous):

@Michele_Laino Still Puzzeled

OpenStudy (michele_laino):

@iam.dhue you are right! please wait a moment...

OpenStudy (michele_laino):

I think that my answer is right,. Please check your question on your textbook, maybe there is a printing error?

OpenStudy (anonymous):

ok Will you continue to help me PLEASE @Michele_Laino

OpenStudy (michele_laino):

your cube root=\[3\sqrt[3]{18}\]

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

I'm ready for other question!